OpenStudy (ttp):

Juanita received two offers to join a video rental program. If she is enrolled in the program for 4 months, which program should she choose? First Program: 6 free video rentals for joining and 1 free video every month Second Program: 2 free video rentals for joining and 2 free videos per month Not enough information The first program The second program Either-She will get the same number of free videos from both

OpenStudy (the_fizicx99):

Second program.

OpenStudy (the_fizicx99):

In a year, she will have 24 + 2 = 26 For the first program she'll have 6 + 12 = 18

OpenStudy (the_fizicx99):

At first yes, she'll have more, however over time the second program will have a greater value

OpenStudy (anonymous):

first program: 6 +1(4)= 10 second program: 2+2(4)=10 is the same for the first 4 months